OpenStudy (nathanjhw):

Sketch a graph of q as a function of p (i.e., q(p)), assuming that the focal length is a constant of 5 centimeters. Show any important features of the graph.

OpenStudy (nathanjhw):

@freckles

OpenStudy (freckles):

Is there no sorta of equation given?

OpenStudy (nathanjhw):

1/p +1/q=1/f

OpenStudy (nathanjhw):

@freckles

OpenStudy (freckles):

oh okay so i think f represents your focal constant so replace f with 5 then solve 1/p+1/q=1/5 for q

OpenStudy (nathanjhw):

p=35/2

OpenStudy (nathanjhw):

q=5p/(p-5)

OpenStudy (freckles):

that bottom this is q written as a function of p with the focal constant as 5

OpenStudy (nathanjhw):

yes

OpenStudy (nathanjhw):

I had to solve this for another part of the question, I just don't know how to graph it now.

OpenStudy (freckles):

\[q=\frac{5p}{p-5}=\frac{5(p-5)+25}{p-5}=\frac{25}{p-5}+5\]

OpenStudy (freckles):

this should tell you the horizontal aysmotote and vertical asymptote?

OpenStudy (freckles):

no ? at the end

OpenStudy (freckles):

lol

OpenStudy (freckles):

you could think about the graph of y=1/x and translate it around to get the graph of y=1/(p-5)+5

OpenStudy (nathanjhw):

So that is how you get the vertical and horizontal asymptotes but what about the rest of the graph?

OpenStudy (freckles):

do you know what y=1/x looks like?

OpenStudy (nathanjhw):

yes

OpenStudy (freckles):